OpenStudy (anonymous):

Find the local maximum and minimum values and saddle point(s) of the function. f(x, y) = x3y + 12x2 - 8y

OpenStudy (anonymous):

i got fx(x,y)= 3x^2y+24 fy(xy)= x^3-8

OpenStudy (anonymous):

fxx (x,y) = 6xy+ 24 fyy(x,y)= 3x^2 so x=0,-8, 2

OpenStudy (anonymous):

What do you mean x=0, -8, 2? What is that the result of?

OpenStudy (anonymous):

setting the fx and fy =0

OpenStudy (anonymous):

i am not sure of what i am suppose to do really past finding the partial derv.

myininaya (myininaya):

fx=3x^2y+24x (this was the only mistake i seen from above so far) fy=x^3-8 fxx=6xy+24 fyy=3x^2

myininaya (myininaya):

i think you meant that since you did fxx right though

OpenStudy (anonymous):

From fx and fy, you should set each to 0 and solve as simultaneous equations and you get (x,y) critical points.

OpenStudy (anonymous):

is fxy= 3x^2?

myininaya (myininaya):

yes

OpenStudy (anonymous):

that is where i got the values for x. those are the critical points, right?

OpenStudy (anonymous):

Critical points are from fx and fy. Set each to 0; solve as simultaneous equations for (x,y) critical points.

OpenStudy (anonymous):

x=0,-8,2

myininaya (myininaya):

fx=0 implies 3x^2y+24x=0 3x(xy+8)=0 3x=0 => x=0 xy+8=0 => x=-8/y fy=0 implies x^3-8=0 when x=2 x=2 => 2=-8/y =>y=-4

OpenStudy (anonymous):

Evaluate fxx at critical points Evaluate fxy at critical points Evaluate fyy at critical points

OpenStudy (anonymous):

Let's call above evaluations A (fxx) B (fxy) C (fyy)

myininaya (myininaya):

it looks that the only only critical number is (2,-4)

myininaya (myininaya):

point*

OpenStudy (anonymous):

If AC-(B)^2>0 and A>0 local min If AC-B^2>0 and A<0 local max If AC-B^2<0 saddle point

OpenStudy (anonymous):

let me work at this

OpenStudy (anonymous):

its the saddle point

OpenStudy (anonymous):

Good work